### v1

@@ 4.3 Results on Decision-based Optimized Image Attack in Image-based AVIs
Table 2  [^4] In-BL., LA-V2, MGPT, m-owl, PGPT, VPGT, OF-2, In-XC., L-1.5, SGPT, Moe represent InstructBLIP, LLaMA-Adapter V2, MiniGPT-4, mPLUG-owl, PandaGPT, VPGTrans, OpenFlamingo-V2, InternLM-XComposer, LLaVA-1.5, ShareGPT4V, Moe-LLaVA, respectively. presents results for decision-based optimized image attacks in image-based AVIs. Regarding visual perception capability, MiniGPT-4 [25] and mPLUG-owl [37] are the most vulnerable LVLMs, while OpenFlamingo-V2 [23] exhibits high robustness (ASR: 21%). For visual knowledge acquisition capability, MiniGPT-4 [25] is the most vulnerable with an ASR of 93%, while InternLM-XComposer [4] performs well with an ASR of 9%. Other LVLMs, including mPLUG-owl [37], OpenFlamingo-V2 [23], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6], have ASRs below 15%, indicating satisfactory performance. In terms of visual reasoning capability, MiniGPT-4 [25] has the highest ASR of 99%, while OpenFlamingo-V2 [23] performs significantly better with an ASR 76% lower than MiniGPT-4. For visual commonsense capability, except for mPLUG-owl [37] (94%), the ASRs for other LVLMs are below 70%, with OpenFlamingo-V2 [23] being the best-performing LVLM. In evaluating object hallucination, OpenFlamingo-V2 [23] remains the top-performing LVLM, while MiniGPT-4 [25] exhibits the poorest performance, achieving a 100% ASR.
@@
To summarize, MiniGPT-4 achieves the highest average ASR at 91%. Other LVLMs with notable performance include OpenFlamingo-V2 [23] at 14%, Moe-LLaVA [6] at 28%, and InternLM-XComposer [4] at 39%. Across various multi-modal capabilities evaluations, the three attack methods consistently demonstrate their effectiveness in the AED evaluation. Combining the PAR attack [40] with the Boundary [41] and Surfree [42] algorithms proves successful in reducing noise amplitude and improving attack detectability. Among the five abilities, visual perception and visual reasoning are the most vulnerable to attacks, with an average ASR of 57%, while visual common sense exhibits the least vulnerability, with an average ASR of 40%. These results emphasize the fragility of LVLMs and serve as a motivation for researchers to develop more robust LVLMs through targeted training approaches. It also highlights the need to enrich defense mechanisms to enhance their resilience against LVLM-agnostic and output probability distribution-agnostic black-box image attacks.
@@
Table 3: Evaluation results of LVLMs’ robustness to text-based AVIs. Cha., Wor., Sen., Sem., Dee.ug, Inp.on represent character-level, word-level, sentence-level, and semantic-level, DeepWordBug, Input-reduction respectively. ASDR ($\downarrow$) is the metric. The best attack category is in bold, the worst is underlined.
@@
Table 4: Evaluation results of LVLMs’ robustness to content bias AVIs. The accuracy ($\uparrow$) is used as the metric. Best results in each row are bold, worst results are underlined. Uns., Cul., Gen. represent unsafe, culture, gender, respectively.
@@
4.4 Results on Black-box Text Attack in Text-based AVIs
Figure 4: Further Analysis: Relationship between robustness to AVIs and (a) Tuning parameters, (b) Vicuna adapters, (c) LLaMA adapters, (d) Training data volume, (e) LLMs. FC, Q--f., Full_Tu., B_Tu., Resa., and RedPajama represent Fully connected layer [36], Q--Former [24], Full Tuning [5], Bias Tuning [9], Resampler [23], and RedPajama--INCITE--Instruct [23] respectively.
Figure 5: The relationship between the LVLMs’ robustness to AVIs and the average score before the attack.
The results of the text-based AVIs are presented in Table 3. Among the different attack methods, TextFooler [49] demonstrated the highest effectiveness with an ASDR of 67%. Conversely, Semantic [11] performed the poorest, achieving an ASDR of only 4%. The low ASDR observed in semantic-level attacks highlights the robustness of LVLMs to instructions provided by individuals with diverse language habits, including Japanese, Chinese, Korean, and others. Among character-level attacks, Pruthi [43] was the most effective, surpassing TextBugger [46] with a 20% higher ASDR. In word-level attacks, TextFooler [49] emerged as the most successful, outperforming BertAttack [48] by 25% in terms of effectiveness. For sentence-level attacks, Input-reduction [45] demonstrated the highest effectiveness, surpassing StressTest [50] by 13%.
@@ 4.4 Results on Black-box Text Attack in Text-based AVIs
4.5 Results on Content Bias AVIs
The experimental results of content bias AVIs are presented in Table 4. Among open-source LVLMs, LLaVA [36] and OpenFlamingo-V2 [23] emerge as the top performers for detecting unsafe information, achieving a 100% detection rate. In contrast, VPGTrans [38] and MiniGPT-4 [25] exhibit lower performance, with detection rates of 14% and 32% respectively. In the context of cultural content bias attacks, LLaVA [36] and OpenFlamingo-V2 [23] continue to demonstrate superior performance, while Otter [3] (46%) and MiniGPT-4 [25] (40%) show poorer results. Regarding racial content bias attacks, BLIP2 [24] emerges as the best-performing model, achieving a detection rate of 90%, while LLaVA [36] performs the worst, with only a 2% detection rate. For gender content bias attacks, ShareGPT4V [5] achieves the highest performance, reaching 89%, while OpenFlamingo-V2 [23] lags with a detection rate of only 9%. Overall, the best performer among all tested open-source LVLMs is ShareGPT4V [5], scoring 74%. VPGTrans [38] performs the worst, with a score of 31%.
Regarding advanced closed-source LVLMs like GeminiProVisionn [1] and GPT-4V [2], while GeminiProVision achieved the top performance among all tested models, we observed that GPT-4V even performed worse than some earlier open-source LVLMs like LLaMA-Adapter V2 [9]. We find that apart from the low detection rate of unsafe information such as hate and self-harm, GPT-4V demonstrates noticeable biases in cultural contexts. For instance, it displays a 25% higher detection rate for American funerals compared to Japanese funerals, and a 10% higher detection rate for American funerals compared to African funerals. We also observed notable instances of racial bias in GeminiProVision, which predicts a higher likelihood of poverty for Black individuals by approximately 30% compared to White individuals. Moreover, significant gender biases are evident as GeminiProVision associates police officers more with males and teachers more with females. These biases hinder the development of fair and reliable LVLMs.
@@ 4.5 Results on Content Bias AVIs
4.6 Further Analysis and Discussion
@@
In this section, we analyze the relationship between the robustness to AVIs and factors such as model structure, training data, and training methods, using evaluation results from various LVLMs. Despite the differences in LVLMs’ structures and training data, the overall framework remains consistent, involving visual encoders, Large Language Models (LLMs), and feature interactors (adapters). While we strive to control variables as much as possible, it is challenging to strictly control them due to variations in models’ configurations. However, this analysis still provides valuable insights and conjectures.
@@
AVIs Robustness and Tuning Parameters. Fig. 4(a) shows that image corruption AVIs exhibit a negative correlation with the quantity of tuning parameters, while content bias AVIs demonstrate a positive correlation. We also find that increasing the tuning parameters from 7B to 7.5B (specifically by tuning the image encoder), enhances the robustness of text-based AVIs and decision-based optimized image AVIs.
@@
AVIs Robustness and LLM Adapters. In Fig.4(b), Vicuna adapters[7] are observed to have minimal impact on text attacks and image corruption. This suggests that relying solely on fully connected layers might pose challenges in effectively mitigating image corruption. On the other hand, Fig.4(c) demonstrates diverse robustness levels in LLaMA[8] adapters, highlighting the importance of prioritizing defense against weaker attack types specific to different adapters.
@@
AVIs and Training Data Volume. In Fig. 4(d), the robustness of LVLMs against different AVIs does not exhibit a significant correlation with the scale of the training data. Instead, we speculate that factors such as data quality, content, and training methods may have a more pronounced impact on LVLMs’ robustness to AVIs.
AVIs and LLMs. In Fig. 4(e), we observe diverse levels of robustness among LVLMs that are based on different LLMs, which highlights the difficulty of adopting a unified defense approach and emphasizes the importance of considering the direction of defense and the specific structural differences across models.
@@
AVIs and Average Score Before Attack. Fig. 5 illustrates the correlation between LVLM’s robustness scores against various attacks and their pre-attack scores. The Pearson Correlation Coefficient, $r$, gauges this correlation. Notably, the relationship with original scores is weaker for image corruptions in image-based AVIs and text-based AVIs, while decision-based optimized black-box attacks in image-based AVIs and content bias AVIs show a stronger correlation, with $r$ values of 0.51 and 0.42, respectively. This suggests that original scores may better predict the robustness of decision-based optimized black-box image-based AVIs and content bias AVIs compared to other attack types. However, the coefficients, $R^{2}$ are low, which means the ability of image and text comprehension may not be well related to the defense against AVIs. We make more detailed analyses considering attack categories and multimodal capabilities in the supplementary material.
5 Conclusion
@@
In conclusion, this paper introduces AVIBench, a comprehensive framework designed to analyze the robustness of Large Vision-Language Models (LVLMs) against different types of adversarial visual-instructions (AVIs), including image-based AVIs, text-based AVIs, and content bias AVIs. AVIBench generates 260K AVIs, encompassing a wide range of multimodal capabilities and content biases. It conducts extensive evaluations involving 14 open-source LVLMs and two closed-source LVLMs. AVIBench provides a valuable tool for assessing the defense mechanisms of LVLMs. The vulnerabilities identified in LVLMs, when subjected to intentional and careless attacks, emphasize the critical need to enhance the robustness, security, and fairness of LVLMs to ensure their responsible deployment across various applications. Additionally, AVIBench will be publicly available as an open-source resource, serving as a foundational tool for robust LVLM research.
@@
ETHICS STATEMENT. This paper analyzes the inherent biases in LVLMs, specifically cultural, gender, and racial biases, and their effectiveness in identifying unsafe information using content bias AVIs. The research aims to promote the safe and fair usage of LVLMs. All images used are sourced from the Internet, and biased images are not intentionally created. The images of content bias attacks will not be publicly shared and will only be used for online testing.

### v2

@@ IV-C Results on Decision-based Optimized Image Attack in Image-based B-AVIs
Table IV presents results for decision-based optimized image attacks in image-based B-AVIs. Regarding visual perception capability, MiniGPT-4 [33] and mPLUG-owl [52] are the most vulnerable LVLMs, while OpenFlamingo-V2 [31] exhibits high robustness (ASR: 21%). For visual knowledge acquisition capability, MiniGPT-4 [33] is the most vulnerable with an ASR of 93%, while InternLM-XComposer [5] performs well with an ASR of 9%. Other LVLMs, including mPLUG-owl [52], OpenFlamingo-V2 [31], LLaVA-1.5 [30], ShareGPT4V [6], and Moe-LLaVA [7], have ASR below 15%, indicating satisfactory performance. In terms of visual reasoning capability, MiniGPT-4 [33] has the highest ASR of 99%, while OpenFlamingo-V2 [31] performs significantly better with an ASR 76% lower than MiniGPT-4. For visual commonsense capability, except for mPLUG-owl [52] (94%), the ASR for other LVLMs are below 70%, with OpenFlamingo-V2 [31] being the best-performing LVLM. In evaluating object hallucination, OpenFlamingo-V2 [31] remains the top-performing LVLM, while MiniGPT-4 [33] exhibits the poorest performance, achieving a 100% ASR.
@@
To summarize, MiniGPT-4 achieves the highest average ASR at 91%. Other LVLMs with notable performance include OpenFlamingo-V2 [31] at 14%, Moe-LLaVA [7] at 28%, and InternLM-XComposer [5] at 39%. Across various multi-modal capabilities evaluations, the three attack methods consistently demonstrate their effectiveness in the AED evaluation. Combining the PAR attack [61] with the Boundary [62] and Surfree [63] algorithms proves successful in reducing noise amplitude and attack detectability. Among the five abilities, visual perception and visual reasoning are the most vulnerable to attacks, with an average ASR of 57%, while visual common sense exhibits the least vulnerability, with an average ASR of 40%. These results emphasize the fragility of LVLMs and serve as a motivation for researchers to develop more robust LVLMs through targeted training approaches. It also highlights the need to enrich defense mechanisms to enhance their resilience against LVLM-agnostic and output probability distribution-agnostic black-box image attacks.
@@
TABLE V: Evaluation results of LVLMs’ robustness to text-based B-AVIs. ASDR ($\downarrow$ indicates the LVLM has greater robustness.) is the metric. The best attack method for each LVLM is in bold, the worst is underlined.
@@
TABLE VI: Evaluation results of LVLMs’ robustness to content bias B-AVIs. The accuracy ($\uparrow$ indicates the LVLM has greater robustness.) is used as the metric. The most robustness LVLM for each content bias is bold, and the worst robustness LVLM for each content bias is underlined. Uns., Cul., Gen. represent unsafe, culture, gender, respectively.
@@
IV-D Results on Black-box Text Attack in Text-based B-AVIs
Figure 5: Further Analysis: Relationship between robustness score to B-AVIs and (a) Tuning parameters, (b) Vicuna adapters, (c) LLaMA adapters, (d) Training data volume, (e) LLMs. FC, Q--f., Full_Tu., B_Tu., Resa., and RedPajama represent Fully connected layer [51], Q--Former [32], Full Tuning [6], Bias Tuning [10], Resampler [31], and RedPajama--INCITE--Instruct [31] respectively.
Figure 6: The relationship between the LVLMs’ robustness score to B-AVIs and the average score before the attack.
The results of the text-based B-AVIs are presented in Table V. Among the different attack methods, TextFooler [67] demonstrated the highest effectiveness with an ASDR of 67%. Conversely, Semantic [18] performed the poorest, achieving an ASDR of only 4%. The low ASDR observed in semantic-level attacks highlights the robustness of LVLMs to instructions provided by individuals with diverse language habits, including Japanese, Chinese, Korean, and others. Among character-level attacks, Pruthi [70] was the most effective, surpassing TextBugger [64] with a 20% higher ASDR. In word-level attacks, TextFooler [67] emerged as the most successful, outperforming BertAttack [66] by 25% in terms of effectiveness. For sentence-level attacks, Input-reduction [72] demonstrated the highest effectiveness, surpassing StressTest [68] by 13%.
@@ IV-D Results on Black-box Text Attack in Text-based B-AVIs
IV-E Results on Content Bias B-AVIs
The experimental results of content bias B-AVIs are presented in Table VI. Among open-source LVLMs, LLaVA [51] and OpenFlamingo-V2 [31] emerge as the top performers for detecting unsafe information, achieving a 100% accuracy. In contrast, VPGTrans [53] and MiniGPT-4 [33] exhibit lower performance, with accuracy of 14% and 32% respectively. In the context of cultural content bias attacks, LLaVA [51] and OpenFlamingo-V2 [31] continue to demonstrate superior performance, while Otter [4] (46%) and MiniGPT-4 [33] (40%) show poorer results. Regarding racial content bias attacks, BLIP2 [32] emerges as the best-performing model, achieving an accuracy of 90%, while LLaVA [51] performs the worst, with only a 2% accuracy. For gender content bias attacks, ShareGPT4V [6] achieves the highest performance, reaching 89%, while OpenFlamingo-V2 [31] lags with an accuracy of only 9%. Overall, the best performer among all tested open-source LVLMs is ShareGPT4V [6], scoring 74%. VPGTrans [53] performs the worst, with a score of 31%.
Regarding advanced closed-source LVLMs like GeminiProVision [2] and GPT-4V [3], while GeminiProVision achieved the top performance among all tested models, we observed that GPT-4V even performed worse than some earlier open-source LVLMs like LLaMA-Adapter V2 [10]. We find that apart from the low accuracy of unsafe information such as hate and self-harm, GPT-4V demonstrates noticeable biases in cultural contexts. For instance, it displays a 25% higher accuracy for American funerals compared to Japanese funerals, and a 10% higher accuracy for American funerals compared to African funerals. We also observed notable instances of racial bias in GeminiProVision, which predicts a higher likelihood of poverty for Black individuals by approximately 30% compared to White individuals. Moreover, significant gender biases are evident as GeminiProVision associates police officers more with males and teachers more with females. These biases hinder the development of fair and reliable LVLMs.
@@ IV-E Results on Content Bias B-AVIs
IV-F Further Analysis and Discussion
@@
In this section, we analyze the relationship between the robustness of B-AVIs and factors such as model structure, training data, and training methods, using evaluation results from various LVLMs. Despite the differences in LVLMs’ structures and training data, the overall framework remains consistent, involving vision encoders, Large Language Models (LLMs), and feature interactors (adapters). While we strive to control variables as much as possible, it is challenging to strictly control them due to variations in models’ configurations. However, this analysis still provides valuable insights and conjectures.
IV-F1 B-AVIs Robustness and Tuning Parameters
@@
In this setting, 3.1M, 28M, 3B, and 7.5B refer to MiniGPT-4 [33], PandaGPT [29], Moe-LLaVA [7], and ShareGPT4V [6]. 107M refers to the average score of VPGTrans [53], BLIP2 [32] and InstructBLIP [28]. 7B refers to the average score of LLaVA [51], InternLM-XComposer [5], and LLaVA-1.5 [30]. Fig. 5(a) shows that image corruption B-AVIs exhibit a negative correlation with the number of tuning parameters, while content bias B-AVIs demonstrate a positive correlation. We also find that increasing the tuning parameters from 7B to 7.5B (specifically by tuning the vision encoder), enhances the robustness of text-based B-AVIs and decision-based optimized image B-AVIs.
IV-F2 B-AVIs Robustness and LLM Adapters
@@
In this setting, w/ FC, w/ LoRA, and Full Tuning refer to MiniGPT-4 [33], PandaGPT [29], and LLaVA [51], respectively. w/ Q--Former refers to the average score of InstructBLIP [28] and VPGTrans [53]. In Fig.5(b), Vicuna adapters[8] are observed to have minimal impact on text attacks and image corruption. This suggests that relying solely on fully connected layers might pose challenges in effectively mitigating image corruption. On the other hand, Fig.5(c) demonstrates diverse robustness levels in LLaMA[9] adapters, highlighting the importance of prioritizing defense against weaker attack types specific to different adapters.
IV-F3 B-AVIs and Training Data Volume
@@
In this setting, 753K, 2.8M, 13.8M, 145M, 204M refer to LLaVA [51], Otter [4], VPGTrans [53], InstructBLIP [28], and mPLUG-owl [52]. In Fig. 5(d), the robustness of LVLMs against different B-AVIs does not exhibit a significant correlation with the scale of the training data. Instead, we speculate that factors such as data quality, content, and training methods may have a more pronounced impact on LVLMs’ robustness to B-AVIs.
IV-F4 B-AVIs and LLMs
In this setting, FlanT5-XL, RedPajama., Qwen, and InternLM refer to BLIP2 [32], OpenFlamingo-V2 [31], Moe-LLaVA [7] and InternLM-XComposer [5], respectively. Vicuna refers to the average score of InstructBLIP [28], LLaVA [51], LLaVA-1.5 [30], MiniGPT-4 [33], PandaGPT [29] and VPGTrans [53]. LLaMA refers to the average score of LLaMA-Adapter V2 [10], mPLUG-owl [52] and Otter [4]. In Fig. 5(e), we observe diverse levels of robustness among LVLMs that are based on different LLMs. The results highlight that it is difficult to adopt a unified defense approach for different LLM-based LVLMs and emphasize the importance of considering both the direction of defense and the specific structural differences across the models.
IV-F5 B-AVIs and Average Score Before Attack
@@
Fig. 6 illustrates the correlation between LVLM’s robustness scores against various attacks and their pre-attack scores. The Pearson Correlation Coefficient, $r$, gauges this correlation. Notably, the relationship with original scores is weaker for image corruptions in image-based B-AVIs and text-based B-AVIs, while decision-based optimized black-box attacks in image-based B-AVIs and content bias B-AVIs show a stronger correlation, with $r$ values of 0.51 and 0.42, respectively. This suggests that original scores may better predict the robustness of decision-based optimized black-box image-based B-AVIs and content bias B-AVIs compared to other attack types. However, the coefficients, $R^{2}$ are low, which means the ability of image and text comprehension may not be well related to the defense against B-AVIs.
V Conclusion
@@
In conclusion, this paper introduces B-AVIBench, a comprehensive framework designed to analyze the robustness of Large Vision-Language Models (LVLMs) against different types of black-box adversarial visual-instructions (B-AVIs), including image-based B-AVIs, text-based B-AVIs, and content bias B-AVIs. B-AVIBench generates 316K B-AVIs, encompassing a wide range of multimodal capabilities and content biases. It conducts extensive evaluations involving 14 open-source LVLMs and two closed-source LVLMs. B-AVIBench provides a valuable tool for assessing the defense mechanisms of LVLMs. The vulnerabilities identified in LVLMs, when subjected to intentional and careless attacks, emphasize the critical need to enhance the robustness, security, and fairness of LVLMs to ensure their responsible deployment across various applications. Additionally, B-AVIBench will be publicly available as an open-source resource, serving as a foundational tool for robust LVLM research.
@@
ETHICS STATEMENT. This paper analyzes the inherent biases in LVLMs. The research aims to promote the safe and fair usage of LVLMs. All images used are sourced from the Internet, and biased images are not intentionally created. The images of content bias attacks will not be publicly shared and will only be used for online testing.